OpenStudy (anonymous):

Find the derivative of e^-x + sinx - cosx I just did a calculus test with a bunch of questions... Many of those that had antiderivative and e really confused me... Please help me to find the answer, but I would also like to know the rules

OpenStudy (anonymous):

the derivative of \(e^x\) is \(e^x\) and similarly the anti derivative of \(e^x\) is also \(e^x\)

OpenStudy (anonymous):

by the chain rule, the derivative of \(e^{f(x)}\) is \(e^{f(x)}f'(x)\)

OpenStudy (anonymous):

for example, the derivative of \(e^{\sin(x)}\) is \(e^{\sin(x)}\cos(x)\)

OpenStudy (anonymous):

and the derivative of \(e^{-x}\) is \(-e^{-x}\)

OpenStudy (anonymous):

Ah... Thank you so much, those are very useful. I was always troubled by e^f(x) and were not so sure before... thanks!